# Percolation in semicontinuum geometries

### v1

@@ Abstract
We study percolation problems of overlapping objects where the underlying geometry is such that in $D-$dimensions, a subset of the directions has a lattice structure, while the remaining directions have a continuum structure. The resulting semicontinuum problem describes the percolation of overlapping shapes in parallel layers or lanes with positional constraints for the placement of the objects along the discrete directions. Several semicontinuum percolation systems are analyzed like hypercuboids with a particular focus on 2D and 3D cases, disks, and parallelograms. Adapting the excluded volume arguments to the semicontinuum setting, we show that for the semicontinuum problem of hypercuboids, for fixed side-lengths of the hypercuboids along the directions in which a lattice structure is maintained, the percolation threshold is always independent of the side-lengths along the continuum directions. The result holds even when there is a distribution for the side-lengths along the continuum directions. Trends in the variation of the thresholds, as we vary the linear measure of the shapes along the continuum directions, are obtained for other semicontinuum models like disks and parallelograms in 2D. The results are compared with those of corresponding continuum and lattice models. For the 2D and 3D models considered, using Monte Carlo simulations, we verify the excluded volume predictions for the trends and numerical values of the percolation thresholds. Very good agreement is seen between the predicted numerical values and the simulation results. The semicontinuum setting also allows us to establish a connection between the percolation problem of overlapping shapes in 2D continuum and triangular lattice. We also verify that the isotropy of the threshold for anisotropic shapes and standard percolation universality class is maintained in the semicontinuum setting.
@@ I Introduction
Over the past several decades, percolation models that exhibit geometric phase transitions have garnered significant attention due to their theoretical and practical importance across various fields  nadiv ; croccolo ; sander ; philcox ; cong ; dechenaux ; pokhrel ; saberi . These models effectively capture the inherent randomness of geometric structures present in many systems and show non-trivial critical behavior isichenko ; adler ; soltani ; sahimi . Traditionally, percolation models have been studied in both lattice and continuum frameworks stauffer ; christensen ; meester . Lattice models are relatively easier to analyze whereas the continuum models are thought to represent better the random geometries encountered in several physical scenarios Lin ; JLin ; thovert ; hunt .
@@
In recent works, researchers have investigated the percolation of overlapping and extended objects or shapes on lattices koza ; koza1 ; brzeski . In these models, objects are said to be overlapping when they share sites. Two objects are deemed connected if they overlap or two of their constituent sites are adjacent to each other. These models serve as a bridge between lattice and continuum approaches to percolation phenomena and find applications in problems such as transport in porous media koponen ; koponen1 . Investigations into the discrete-to-continuum transition of overlapping aligned squares and cubes on a lattice have been conducted in koza ; koza1 . A similar study for hyperspheres was carried out in brzeski .
@@
In recent research, we demonstrated that models involving extended, aligned, and overlapping objects on lattices exhibit unique properties compared to their continuum counterparts jasna . Notably, these models display threshold behavior that depends on the width of the objects, which contrasts with the aspect-ratio-independent behavior observed in corresponding continuum models. For example, it was found that for overlapping and aligned rectangles placed on a square lattice, the percolation threshold decreases with increasing length when the width is one, increases with length when the width is greater than two, and remains constant when the width is two jasna .
@@
A lattice percolation model that involves extended overlapping objects is expected to transition to its corresponding continuum model by taking the limit where the lattice spacing approaches zero in all directions while proportionally increasing the size of the objects koza ; koza1 . In Ref. jasna , while discussing the percolation of extended shapes on lattices, we proposed the possibility of taking the continuum limit of such lattice models only along a subset of the possible directions. In this work, we explore this scenario in detail, where, in a percolation model, the lattice structure is retained in one or more directions, while a continuum structure is assumed along the remaining directions. We call the resulting percolation model a semicontinuum one. Percolation models of overlapping objects in such semicontinuum geometries and their critical thresholds are the focus of the present work.
@@
In simple terms, the semicontinuum model we consider here can be described as a percolation model involving overlapping shapes in parallel layers or lanes (see Fig. 1). Within each layer, the shapes can freely overlap. However, their placement and connectivity across layers are constrained by the discrete geometry present in that direction. Such a layered structure bears some resemblance to many real-world formations. Situations one can think of include materials with a layered structure, including rock sediments bai ; xu or many biological tissues exhibiting layered structures bini ; hong . The random geometries found in such layered formations may be effectively captured by semicontinuum geometries. Beyond their potential suitability to describe random structures in such settings, these models occupy an intermediate space between lattice and continuum geometries. As we discuss in this paper, they can exhibit several non-trivial properties not observed in purely lattice or continuum systems. Understanding the semicontinuum models can therefore provide deeper insights into the role of discrete and continuum geometries in determining the physical properties of systems and how these properties evolve as we transition from discrete to continuum spaces.
@@
In this paper, we investigate the percolation properties of semicontinuum systems composed of overlapping objects. Specifically, we examine hypercuboids in general $D$ dimensions, disks in 2D, and parallelograms in 2D. The specific cases of hypercuboids in physically significant dimensions of 2D and 3D, i.e, that of rectangles and cuboids are considered in more detail. Making use of the approximate analytical method of Excluded Volume argument balberg ; balberg1 adapted to the semicontinuum setting, we show that for fixed side-lengths of the hypercuboids along the directions in which a lattice structure is maintained, the percolation thresholds are found to be independent of the side-lengths along the continuum directions. This remains true even if there is a distribution for the side-lengths along the continuum directions. This behavior in the semicontinuum model is notably different from what is seen in both lattice and continuum systems. For example in the 2D lattice model, the trend of the threshold as the length of the rectangle varies depends on its width (it may increase, decrease, or remain constant, as discussed in Ref. jasna ), whereas in continuum models, the threshold is always a fixed constant. For the semicontinuum model of disks in 2D, the percolation threshold is found to be a monotonically decreasing function of radius in contrast to the radius-independent threshold in continuum  meester ; balram . Examining the semicontinuum model of parallelograms in 2D allows us to establish a connection between the percolation problem of overlapping shapes in 2D continuum space and the triangular lattice.
@@
The paper is structured as follows. In Sec. II, we precisely define the semicontinuum model of percolation. In Sec. III, using a version of the excluded volume argument adapted to the semicontinuum geometry, we obtain numerical estimates and trends in the thresholds for several overlapping shapes in 2D and 3D, and analyze the case of hypercuboids in higher dimensions. Predictions in 2D and 3D are compared with simulation results. The impact of discreteness in a subset of directions upon percolation threshold is discussed in detail. Finally, we conclude the paper in Sec. IV.
II Model definition
@@
As we increase the number of rectangles, the system undergoes a percolation transition. In a finite system, a spanning cluster appears for the first time at the critical point, marking the percolation transition. We can characterize the critical point or the percolation threshold in terms of the areal density $\eta$ or the covered area fraction $\phi$ as we do in the usual continuum percolation problems involving overlapping objects mertens ; xia ; baker . The areal density $\eta$ refers to the average total area of all objects placed in a unit area of the plane, while the covered area fraction $\phi$ represents the average fraction of the plane occupied by the objects. They are related by mertens ,
@@
We define a spanning cluster as a connected path of objects that spans the system. In finite systems, we will consider spanning in the horizontal and vertical directions separately, allowing us to define horizontal and vertical percolation thresholds. However, as we will see, in the thermodynamic limit, both thresholds converge to a single value, similar to what is observed in lattice and continuum percolation problems involving anisotropic objects jasna ; klatt . To minimize finite-size effects, periodic boundary conditions are assumed in both the directions. Typical configurations showing non-percolating and percolating phases in the 2D problem defined are shown in Fig. 2
@@
III Results and Discussion
III.1 Excluded volume theory for the semicontinuum models
The excluded volume argument has traditionally been widely used to analyze continuum percolation systems balberg ; balberg1 ; alvarez ; balberg2 . Let $V_{ex}$ represent the volume of the region around an object within which if another object is placed, they will overlap. The fundamental tenet of the excluded volume argument is that at the critical point, the total excluded volume $B_{c}$ which is the product of the number density $n_{c}$ of objects and their excluded volume $V_{ex}$ remains constant for similar shapes in a given dimension balberg ; balberg1 . I.e, in the relation,
@@ III.1 Excluded volume theory for the semicontinuum models
$B_{c}$ is expected to be a constant for continuum systems composed of shapes that are related to each other by scaling and have the same orientation klatt in a given dimension (for example, disks and aligned ellipses). Since the number density of objects is given by $n=\eta/V$, where $V$ is the volume of an object, using Eq. (1) and Eq.(2), the critical covered volume fraction (ccvf) is given by,
@@
Recently, Koza et.al extended the excluded volume argument to study the percolation of extended shapes on lattices koza . In that context, the excluded volume is represented by a connectedness factor which is defined as the number of ways in which two basic percolating units can connect on the lattice koza .
@@
In fact, this is the expression for ccvf in the limit of $k_{1}\rightarrow\infty$ of the lattice model of rectangles discussed in Ref. jasna . We can see from Eq. (4) that for the semicontinuum system, for a fixed width $k_{2}$, the percolation threshold $\phi_{c}^{k_{1},k_{2}}$ is a constant, regardless of length $k_{1}$. For different widths, we get different constant thresholds which are independent of the length. This behavior contrasts sharply with the typical continuum scenario, where the percolation threshold is independent of the aspect ratio of the rectangles klatt . Let us consider the specific case of squares in a 2D square lattice ($k_{1}=k_{2}=k$) and compare it with the case of squares in a semicontinuum geometry. As $k\rightarrow\infty$, the threshold in both cases will approach the continuum value. From Eq. (4) with $k_{2}=k$, we observe that in the semicontinuum geometry, $\phi_{c}^{k}$ increases monotonically with $k$. This contrasts with the non-monotonic behavior of the threshold observed for overlapping squares on a lattice koza .
If we take a lattice model with rectangles of size $k_{1}\times k_{2}$, and take the continuum limit $k_{1}\rightarrow\infty$ keeping $k_{2}$ finite, we essentially obtain the semicontinuum model (See Ref. jasna ). The approach of the lattice thresholds to the semicontinuum thresholds for different values of $k_{2}$ was considered in Ref. jasna and is depicted in Fig. 5. The numerical value of $B_{c}$ is assumed to be the same as that of the corresponding continuum system  koza . We can see that, for widths $k_{2}=1$ and $k_{2}=3$, the thresholds in the lattice model approach the semicontinuum thresholds with opposing trends. However, for $k_{2}=2$, the lattice and semicontinuum thresholds coincide and are independent of $k_{1}$. The case of overlapping rectangles on purely lattice systems is discussed in detail in Ref. jasna .
@@
From Eq. (9), we can conclude that for fixed values of $k_{2}$ and $k_{3}$, the percolation threshold is independent of $k_{1}$.
@@
$V_{ex}=\left[\sum_{\begin{subarray}{c}j_{1},j_{2},...j_{d}=1\\ l_{1},l_{2},...l_{d}=1\end{subarray}}^{M}p_{1}^{j_{1}}p_{2}^{j_{2}}...p_{d}^{j_{d}}p_{1}^{l_{1}}p_{2}^{l_{2}}...p_{d}^{l_{d}}\prod_{i=1}^{d}\left(k_{i}^{j_{i}}+k_{i}^{l_{i}}\right)\right]$
×$\times\left[\prod_{i=d+1}^{D}(2k_{i}-1)\right]\left[1+2\sum_{i=d+1}^{D}\dfrac{1}{(2k_{i}-1)}\right]$ (11)
@@
$V_{ex}=2^{d}\left[\sum_{j_{1},j_{2},...j_{d}=1}^{M}p_{1}^{j_{1}}p_{2}^{j_{2}}...p_{d}^{j_{d}}k_{1}^{j_{1}}k_{2}^{j_{2}}...k_{d}^{j_{d}}\right]$
×$\times\left[\prod_{i=d+1}^{D}(2k_{i}-1)\right]\left[1+2\sum_{i=d+1}^{D}\dfrac{1}{(2k_{i}-1)}\right]$ (13)
@@
$\phi_{c}^{k_{d+1},k_{d+2},...,k_{D}}=$
$1-\exp\left(-B_{c}\frac{\left[\displaystyle\prod_{i=d+1}^{D}k_{i}\right]}{2^{d}\left[\displaystyle\prod_{i=d+1}^{D}(2k_{i}-1)\right]\left[1+2\displaystyle\sum_{i=d+1}^{D}\dfrac{1}{(2k_{i}-1)}\right]}\right)$ (14)
@@
Figure 5: Comparison of excluded volume predictions of percolation thresholds for overlapping rectangles on a lattice and overlapping rectangles on lanes (semicontinuum model) for different widths $k_{2}$. The continuum threshold of rectangles is also shown ($\phi_{c}\approx 0.667$klatt ) for comparison.
@@
In the following section, we present the simulation results for the semicontinuum models of rectangles and disks in 2D, and cuboids in 3D and verify the above results. Additionally, we discuss the case of parallelograms in lanes and its relation to the percolation of overlapping parallelograms on a 2D triangular lattice jasna .
III.2 Simulation Results
III.2.1 Squares and Rectangles in 2D
We first consider the $2D$ semicontinuum percolation problem of overlapping squares and rectangles. The objects are randomly placed in an $L\times L$ system one by one until a spanning cluster appears for the first time in vertical and horizontal directions. The process is repeated multiple times to determine the horizontal and vertical percolation probabilities as functions of the covered volume fraction. This is done by determining the fraction of realizations in which the respective type of spanning cluster appears for the first time. System sizes considered range from a minimum of $L=48$ to a maximum of $L=512$ and a minimum of $10^{3}$ realizations are considered in all cases. Typical plots of percolation probability against covered area fraction $\phi$ obtained for different system sizes are shown in Fig. 7 (a).
Figure 7: (a) Percolation probability vs covered area fraction $\phi$ for rectangles in lanes with $k_{1}=2$, $k_{2}=1$. Vertical percolation probability is plotted. (b) Data collapse for the semicontinuum system of $2\times 1$ rectangles. Percolation probability is plotted against $(\phi-\phi_{c})L^{1/\nu}$ for different $L$ with $\phi_{c}\approx 0.531$ and $0.751/\nu=$ 0.75. Curves for different $L$ fall on top of each other.
From the percolation probability curves, we evaluate the effective percolation threshold $\phi_{c}(L)$ for different finite system sizes $L$, as well as the width of the transition region $\Delta(L)$, by fitting the curves with an error function $\frac{1+erf[(\phi-\phi_{c}(L))/\Delta(L)]}{2}$ rintoul . To calculate the actual threshold in the thermodynamic limit $\phi_{c}(\infty)$, we plot $\phi_{c}(L)$ against $\Delta(L)$ which obey the scaling relation stauffer ,

### v2

@@ Abstract
We study percolation problems of overlapping objects where the underlying geometry is such that in $D-$dimensions, a subset of the directions has a lattice structure, while the remaining directions have a continuum structure. The resulting semicontinuum problem describes the percolation of overlapping shapes in parallel layers or lanes with positional constraints for the placement of the objects along the discrete directions. Several semicontinuum percolation systems are analyzed like hypercuboids with a particular focus on 2D and 3D cases, disks, and parallelograms. Adapting the excluded volume arguments to the semicontinuum setting, we show that for the semicontinuum problem of hypercuboids, for fixed side-lengths of the hypercuboids along the directions in which a lattice structure is maintained, the percolation threshold is always independent of the side-lengths along the continuum directions. The result holds even when there is a distribution for the side-lengths along the continuum directions. Trends in the variation of the thresholds, as we vary the linear measure of the shapes along the continuum directions, are obtained for other semicontinuum models like disks and parallelograms in 2D. The results are compared with those of corresponding continuum and lattice models. For the 2D and 3D models considered, using Monte Carlo simulations, we verify the excluded volume predictions for the trends and numerical values of the percolation thresholds. Very good agreement is seen between the predicted numerical values and the simulation results. The semicontinuum setting also allows us to establish a connection between the percolation problem of overlapping shapes in 2D continuum and the triangular lattice. We also verify that the isotropy of the threshold for anisotropic shapes and standard percolation universality class is maintained in the semicontinuum setting.
@@ 1 Introduction
Over the past several decades, percolation models that exhibit geometric phase transitions have garnered significant attention due to their theoretical and practical importance across various fields  [1, 2, 3, 4, 5, 6, 7, 8]. These models effectively capture the inherent randomness of geometric structures present in many systems and show non-trivial critical behavior [9, 10, 11, 12]. Traditionally, percolation models have been studied in both lattice and continuum frameworks [13, 14, 15]. Lattice models are relatively easier to analyze whereas the continuum models are thought to represent better the random geometries encountered in several physical scenarios [16, 17, 18, 19].
@@
In recent works, researchers have investigated the percolation of overlapping and extended objects or shapes on lattices [20, 21, 22]. In these models, objects are said to be overlapping when they share sites. Two objects are deemed connected if they overlap or two of their constituent sites are adjacent to each other. These models serve as a bridge between lattice and continuum approaches to percolation phenomena and find applications in problems such as transport in porous media [23, 24]. Investigations into the discrete-to-continuum transition of overlapping aligned squares and cubes on a lattice have been conducted in [20, 21]. A similar study for hyperspheres was carried out in [22].
@@
In recent research, we demonstrated that models involving extended, aligned, and overlapping objects on lattices exhibit unique properties compared to their continuum counterparts [25]. Notably, these models display threshold behavior that depends on the width of the objects, which contrasts with the aspect-ratio-independent behavior observed in corresponding continuum models. For example, it was found that for overlapping and aligned rectangles placed on a square lattice, the percolation threshold decreases with increasing length when the width is one, increases with length when the width is greater than two, and remains constant when the width is two [25].
@@
A lattice percolation model that involves extended overlapping objects is expected to transition to its corresponding continuum model by taking the limit where the lattice spacing approaches zero in all directions while proportionally increasing the size of the objects [20, 21]. In Ref. [25], while discussing the percolation of extended shapes on lattices, we proposed the possibility of taking the continuum limit of such lattice models only along a subset of the possible directions. In this work, we explore this scenario in detail, where, in a percolation model, the lattice structure is retained in one or more directions, while a continuum structure is assumed along the remaining directions. We call the resulting percolation model a semicontinuum one. Percolation models of overlapping objects in such semicontinuum geometries and their critical thresholds are the focus of the present work.
@@
In simple terms, the semicontinuum model we consider here can be described as a percolation model involving overlapping shapes in parallel layers or lanes (see Fig. 1). Within each layer, the shapes can freely overlap. However, their placement and connectivity across layers are constrained by the discrete geometry present in that direction. Such a layered structure bears some resemblance to many real-world formations. Situations one can think of include materials with a layered structure, including rock sediments [26, 27] or many biological tissues exhibiting layered structures [28, 29]. The random geometries found in such layered formations may be effectively captured by semicontinuum geometries. Beyond their potential suitability to describe random structures in such settings, these models occupy an intermediate space between lattice and continuum geometries. As we discuss in this paper, they can exhibit several non-trivial properties not observed in purely lattice or continuum systems. Understanding the semicontinuum models can therefore provide deeper insights into the role of discrete and continuum geometries in determining the physical properties of systems and how these properties evolve as we transition from discrete to continuum spaces.
@@
In this paper, we investigate the percolation properties of semicontinuum systems composed of overlapping objects. Specifically, we examine hypercuboids in general $D$ dimensions, disks in 2D, and parallelograms in 2D. The specific cases of hypercuboids in physically significant dimensions of 2D and 3D, i.e, that of rectangles and cuboids are considered in more detail. Making use of the approximate analytical method of Excluded Volume argument [30, 31] adapted to the semicontinuum setting, we show that for fixed side-lengths of the hypercuboids along the directions in which a lattice structure is maintained, the percolation thresholds are found to be independent of the side-lengths along the continuum directions. This remains true even if there is a distribution for the side-lengths along the continuum directions. This behavior in the semicontinuum model is notably different from what is seen in both lattice and continuum systems. For example in the 2D lattice model, the trend of the threshold as the length of the rectangle varies depends on its width (it may increase, decrease, or remain constant, as discussed in Ref. [25]), whereas in continuum models, the threshold is always a fixed constant. For the semicontinuum model of disks in 2D, the percolation threshold is found to be a monotonically decreasing function of radius in contrast to the radius-independent threshold in continuum  [15, 32]. Examining the semicontinuum model of parallelograms in 2D allows us to establish a connection between the percolation problem of overlapping shapes in 2D continuum space and the triangular lattice.
@@
The paper is structured as follows. In Sec. 2, we precisely define the semicontinuum model of percolation. In Sec. 3, using a version of the excluded volume argument adapted to the semicontinuum geometry, we obtain numerical estimates and trends in the thresholds for several overlapping shapes in 2D and 3D, and analyze the case of hypercuboids in higher dimensions. Predictions in 2D and 3D are compared with simulation results. The impact of discreteness in a subset of directions upon percolation threshold is discussed in detail. Finally, we conclude the paper in Sec. 4.
2 Model definition
@@
As we increase the number of rectangles, the system undergoes a percolation transition. In a finite system, a spanning cluster appears for the first time at the critical point, marking the percolation transition. We can characterize the critical point or the percolation threshold in terms of the areal density $\eta$ or the covered area fraction $\phi$ as we do in the usual continuum percolation problems involving overlapping objects [33, 34, 35]. The areal density $\eta$ refers to the average total area of all objects placed in a unit area of the plane, while the covered area fraction $\phi$ represents the average fraction of the plane occupied by the objects. They are related by [33],
@@
We define a spanning cluster as a connected path of objects that spans the system. In finite systems, we will consider spanning in the horizontal and vertical directions separately, allowing us to define horizontal and vertical percolation thresholds. However, as we will see, in the thermodynamic limit, both thresholds converge to a single value, similar to what is observed in lattice and continuum percolation problems involving anisotropic objects [25, 36]. To minimize finite-size effects, periodic boundary conditions are assumed in both the directions. Note that the “spanning”  criteria with periodic boundary condition is the same as considering an arbitrary cut in the periodic system and checking for spanning. This has been used in some of the earlier works (See for example [37]) and is expected to give the same threshold values in the infinite system limit as other criterion that can be used to define an effective threshold for finite systems. Typical configurations showing non-percolating and percolating phases in the 2D problem defined are shown in Fig. 2
@@
3 Results and Discussion
3.1 Excluded volume theory for the semicontinuum models
The excluded volume argument has traditionally been widely used to analyze continuum percolation systems [30, 31, 38, 39]. Let $V_{ex}$ represent the volume of the region around an object within which if another object is placed, they will overlap. The fundamental tenet of the excluded volume argument is that at the critical point, the total excluded volume $B_{c}$ which is the product of the number density $n_{c}$ of objects and their excluded volume $V_{ex}$ remains constant for similar shapes in a given dimension [30, 31]. I.e, in the relation,
@@ 3.1 Excluded volume theory for the semicontinuum models
$B_{c}$ is expected to be a constant for continuum systems composed of shapes that are related to each other by scaling and have the same orientation [36] in a given dimension (for example, disks and aligned ellipses). Since the number density of objects is given by $n=\eta/V$, where $V$ is the volume of an object, using Eq. (1) and Eq.(2), the critical covered volume fraction (ccvf) is given by,
@@
Recently, Koza et al. extended the excluded volume argument to study the percolation of extended shapes on lattices [20]. In that context, the excluded volume is represented by a connectedness factor which is defined as the number of ways in which two basic percolating units can connect on the lattice [20].
@@
In fact, this is the expression for ccvf in the limit of $k_{1}\rightarrow\infty$ of the lattice model of rectangles discussed in Ref. [25]. We can see from Eq. (4) that for the semicontinuum system, for a fixed width $k_{2}$, the percolation threshold $\phi_{c}^{k_{1},k_{2}}$ is a constant, regardless of length $k_{1}$. For different widths, we get different constant thresholds which are independent of the length. This behavior contrasts sharply with the typical continuum scenario, where the percolation threshold is independent of the aspect ratio of the rectangles [36]. Let us consider the specific case of squares in a 2D square lattice ($k_{1}=k_{2}=k$) and compare it with the case of squares in a semicontinuum geometry. As $k\rightarrow\infty$, the threshold in both cases will approach the continuum value. From Eq. (4) with $k_{2}=k$, we observe that in the semicontinuum geometry, $\phi_{c}^{k}$ increases monotonically with $k$. This contrasts with the non-monotonic behavior of the threshold observed for overlapping squares on a lattice [20].
If we take a lattice model with rectangles of size $k_{1}\times k_{2}$, and take the continuum limit $k_{1}\rightarrow\infty$ keeping $k_{2}$ finite, we essentially obtain the semicontinuum model (See Ref. [25]). The approach of the lattice thresholds to the semicontinuum thresholds for different values of $k_{2}$ was considered in Ref. [25] and is depicted in Fig. 5. The numerical value of $B_{c}$ is assumed to be the same as that of the corresponding continuum system  [20]. We can see that, for widths $k_{2}=1$ and $k_{2}=3$, the thresholds in the lattice model approach the semicontinuum thresholds with opposing trends. However, for $k_{2}=2$, the lattice and semicontinuum thresholds coincide and are independent of $k_{1}$. The case of overlapping rectangles on purely lattice systems is discussed in detail in Ref. [25].
@@
From Eq. (9), we can conclude that for fixed values of $k_{2}$ and $k_{3}$, the percolation threshold is independent of $k_{1}$. In fact such independence of thresholds on the side-length(s) or measure(s) of the shapes along the continuum direction(s) in the model is expected because uniformly stretching the system along the continuum direction(s) maintains its connectivity structure. The same fact is reflected in the cancellation of factors corresponding to the continuum direction(s) in the ratio of volume to excluded volume of the objects in the excluded volume arguments leading to Eq. (4), Eq. (8), and Eq. (9).
@@
$V_{ex}=\left[\sum_{\begin{subarray}{c}j_{1},j_{2},...j_{d}=1\\ l_{1},l_{2},...l_{d}=1\end{subarray}}^{M}p_{1}^{j_{1}}p_{2}^{j_{2}}...p_{d}^{j_{d}}p_{1}^{l_{1}}p_{2}^{l_{2}}...p_{d}^{l_{d}}\prod_{i=1}^{d}\left(k_{i}^{j_{i}}+k_{i}^{l_{i}}\right)\right]\times\left[\prod_{i=d+1}^{D}(2k_{i}-1)\right]\left[1+2\sum_{i=d+1}^{D}\dfrac{1}{(2k_{i}-1)}\right]$ (11)
@@
$V_{ex}=2^{d}\left[\sum_{j_{1},j_{2},...j_{d}=1}^{M}p_{1}^{j_{1}}p_{2}^{j_{2}}...p_{d}^{j_{d}}k_{1}^{j_{1}}k_{2}^{j_{2}}...k_{d}^{j_{d}}\right]\times\left[\prod_{i=d+1}^{D}(2k_{i}-1)\right]\left[1+2\sum_{i=d+1}^{D}\dfrac{1}{(2k_{i}-1)}\right]$ (13)
@@
$\phi_{c}^{k_{d+1},k_{d+2},...,k_{D}}=1-\exp\left(-B_{c}\frac{\left[\displaystyle\prod_{i=d+1}^{D}k_{i}\right]}{2^{d}\left[\displaystyle\prod_{i=d+1}^{D}(2k_{i}-1)\right]\left[1+2\displaystyle\sum_{i=d+1}^{D}\dfrac{1}{(2k_{i}-1)}\right]}\right)$ (14)
@@
Figure 5: Comparison of excluded volume predictions of percolation thresholds for overlapping rectangles on a lattice and overlapping rectangles on lanes (semicontinuum model) for different widths $k_{2}$. The continuum threshold of rectangles is also shown ($\phi_{c}\approx 0.667$[36]) for comparison.
@@
In the following section, we present the simulation results for the semicontinuum models of rectangles and disks in 2D, and cuboids in 3D and verify the above results. Additionally, we discuss the case of parallelograms in lanes and its relation to the percolation of overlapping parallelograms on a 2D triangular lattice [25].
3.2 Simulation Results
3.2.1 Squares and Rectangles in 2D:
We first consider the $2D$ semicontinuum percolation problem of overlapping squares and rectangles. The objects are randomly placed in an $L\times L$ system one by one until a spanning cluster appears for the first time in vertical and horizontal directions. The process is repeated multiple times to determine the horizontal and vertical percolation probabilities as functions of the covered volume fraction. This is done by determining the fraction of realizations in which the respective type of spanning cluster appears for the first time. We use the Newman-Ziff algorithm for determining the thresholds [40] utilizing a grid-based union-find algorithm for the semicontinuum geometry which is similar to the algorithm used in Ref. [33] for continuum systems. We divide the $L\times L$ semicontinuum system into several grids of size $k_{1}\times k_{2}$. We ensure that $L$ is always an integer multiple of $k_{1}$ in the $X$ direction and a multiple of $k_{2}$ in the $Y$ direction which simplifies the process of checking for overlapping. System sizes considered range from a minimum of $L=32$ to a maximum of $L=4096$ ($L=8190$ for larger $k$ squares) and the number of realizations considered ranges from a minimum of $10^{4}$ to a maximum of $10^{6}$. Typical plots of percolation probability against covered area fraction $\phi$ obtained for a few different system sizes are shown in Fig. 7 (a).
Figure 7: (a) Percolation probability vs covered area fraction $\phi$ for rectangles in lanes with $k_{1}=2$, $k_{2}=1$. Vertical percolation probability is plotted. (b) Data collapse for the semicontinuum system of $2\times 1$ rectangles. Percolation probability is plotted against $(\phi-\phi_{c})L^{1/\nu}$ for different $L$ with $\phi_{c}\approx 0.5311$ and $0.751/\nu=$ 0.75. Curves for different $L$ fall on top of each other.
From the percolation probability curves, we evaluate the effective percolation threshold $\phi_{c}(L)$ for different finite system sizes $L$, as well as the width of the transition region $\Delta(L)$, by fitting the curves with an error function $\frac{1+erf[(\phi-\phi_{c}(L))/\Delta(L)]}{2}$ [41]. To calculate the actual threshold in the thermodynamic limit $\phi_{c}(\infty)$, we plot $\phi_{c}(L)$ against $\Delta(L)$ which obey the scaling relation [13],